# Sequential Quantum Measurements and the Instrumental Group Algebra

### v1

@@ III.3 Sequential Measurements: Kraus-Operator Distributions and the Convolution (2 of 3)
Equations 123 and 125 are deeper elaborations of equations 87 and 85.
IV Kraus-Operator Densities, Ultraoperators, and the IG Algebra
@@
Just like the real line or the complex plane, the IG is the domain of an entire suite of analytic tools — with invariant integrals (IV.1), invariant derivatives (IV.2), involutive symmetries (IV.3), and intertwined differential and algebraic equations (IV.4)— where functions such as the Kraus-operator density (KOD) can be calculated, propogated, and varied.
IV.1 The Haar Measures: Kraus-Operator Densities and the Convolution (3 of 3)
@@
As seen in sections III.1 and III.2 and to be explained further in section IV.4, the samples of a weak-instrument Kraus-operator distribution are the increments of a continuous invariant Markov process, increments that are locally native to another space beside the IG, or rather tangent to it. In particular, this means that it is usually easier for small times to work directly with the Kraus-operator distribution of the weak instrument and not the KOD. Therefore, it is useful to notice that the KOD of a convolution can be computed by directly applying the Kraus-operator distribution of one instrument to the KOD of the other. There are two such expressions in this regard, one for the KODs with respect to the left-invariant measure,
@@
IV.2 Translation Ultraoperators and their Invariant Derivatives
@@
IV.3 The IG Algebra and Two Involutions: Gelfand-Kolmogorov and Cartan-Schmidt
@@
There are a couple of important subspaces of the IGA that come up. Although here is not the place to dig into them too much, they are mentioned because the KODs which solve the Kolmogorov equation to be introduced in section IV.4 are indeed in these subspaces for finite times. The first important subspace is the closure of the space of infinitely-differentiable functions with compact support, sometimes called the Frechet space [74]. The dual space of this Frechet space is called the space of distributions. The term “distribution” here refers to “functions” like Dirac delta-functions and their derivatives which although related are not to be confused with the Kraus-operator distribution, where the KOD is actually in the Frechet space, on the other side of the function duality from the “distributions”. The second important subspace is the so-called Schwartz space for semisimple IGs, their dual space known as the tempered distributions. [75] If the KOD is in the Schwartz space, then the corresponding instrument-element representations are examples of the so-called tempered representations. [76] However, for quantum systems the instrument-element representation is generally not a unitary representation as will be explained in just a moment.
@@
which will be called the Kolmogorov (ultraoperator) adjoint because of its application in section IV.4. The use of an , called the “Ankh” or “key to life”, is mostly because every obvious choice to denote an adjoint has already been taken — * for convolution, $\dagger$} for Hermitian conjugation, $\ddagger$} for the Hilbert-Schmidt adjoint (equation 75) , and $\ddagger$ for the Choi-Jamiolkowski quasi-adjoint (equation 77). As far as the author is aware, the Ankh does not have any political or historical associations that might make it a distastful symbol to use. If anything, the Ankh has a fun turn-of-the-twentieth century Western alchemical vibe.
@@
Meanwhile, the right-translation ultraoperators are not stochastic unless the IG is unimodular. We will have no further use for the right-translation ultraoperators, so this point will not be discussed further. Remember that the integration has been chosen with respect to the left-invariant Haar measure because the time of the measuring process is always considered to march from right to left, as explained at the end of section IV.1.
@@
where in particular $d_{\mathrm{L}}(x^{\dagger})=d_{\mathrm{R}}x\,$ because the right-invariant measure is unique [66, p. 26]. In the same way as was explained for the , the IGA is not quite a $\text{C}^{\ddagger}$-algebra” either, but I believe it too defines another enveloping C*-algebra by the same standard construction [30, p. 47]. The enveloping $\text{C}^{\ddagger}$-algebra is clearly relevant to continuous quantum measurement. Nevertheless, it is a technical perspective that does not seem to appear yet in quantum measurement circles.
@@
IV.4 The Kolmogorov Equation for the KOD of a Continuous Measurement
Let $\mathcal{I}_{dt}$ be a weak measurement such as those found in sections III.1 and III.2 and $d\mu_{T}(x)$ be the Kraus-operator distribution of a continuous instrument
@@ IV.4 The Kolmogorov Equation for the KOD of a Continuous Measurement
Meanwhile, let the KOD — introduced in section IV.1 (see equations 130 and 147) — of the instrument at time $T$ be denoted as
@@
as explained around equations 89 and 85, but so do the Kraus-operator distribution and its KOD:
@@
which can be thought of as special cases of equations 123, 125, and 136. Equations such as 213 are examples of so-called Chapman-Kolmogorov equations.[3, 79, 80, 81]
Let $\mathcal{O}_{x}$ be an instrument-element representation of the IG (as explained in section III.3.) Observe in particular that they intertwine the left-translation ultraoperators (introduced in section IV.2) with left composition by the corresponding superoperator from the instrument-element representation,
@@
it is clear that the total operation superoperator $\mathcal{Z}_{dt}$ acting on the left of the instrument-element representation $\mathcal{O}_{x}$ by composition is equivalent to the left-convolution ultraoperator (also introduced in section IV.2)
@@
where is the Kolmogorov adjoint over the IG (introduced in section IV.3.) Meanwhile, the adjoint of the total operation ultraoperator of the weak instrument will have an infinitesimal generator,
@@
as introduced in section III.1 (equations 90 and 92.) Simultaneous measurements can also be considered, but for simplicity only one Lindblad operator will be considered at first. These weak instruments define IGs (also explained in section III.1.) Let the instrument elements of these weak instruments be “minimally” separated into a Kraus-operator distribution and an instrument-element representation in the following way:
@@ V Concluding Remarks
In addition to the Kolmogrov generator, further operators on the IGA, generally called ultraoperators, were also considered, allowing the analysis to be completely autonomous from the superoperators that originally represented the measuring process. Perhaps the most fundamental of the ultraoperators were the right-invariant derivatives and left-translations, denoted $\underleftarrow{X}$ and $\mathcal{L}_{x}$. With these, the left-convolution ultraoperators could be constructed in direct analogy with the quantum channel superoperators,
@@
Table 5: Ultraoperator-Superoperator Intertwining Relations: The instrument-element representations of the IG, $\mathcal{O}_{x}^{(\lambda)}$ define an intertwiner between the algebra of ultraoperators (acting on the IGA) and the algebra of superoperators (acting on the dual of the von Neumann algebra). This table displays four pairs of operators, intertwined by the instrument-element representation, that are particularly relevant to continuous measurements and sequential measurements. After the IGA and ultraoperator are introduced, these intertwining relations were derived and used in section §IV.4. Concepts with the ※  mark were first introduced in previous work [3].
@@
Having said all of this, it is crucial to return to the other primary tool of the IMP that was virtually abandoned at the beginning of this paper, the so-called modified Maurer-Cartan stochastic differential or MMCSD. The MMCSD is precisely what an automated measuring instrument would use to update its IG position as it continues to observe. Further, such automation is only possible if the IG is automatable — that is, itself “automatic” (section III.1).
@@ Acknowledgements.
Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities.
Appendix A System-Meter-Interaction Models
@@ A.2 Diffusive Processes
Appendix B Invariant One-Forms, Haar Measures, and the Modular Function
@@
Some further properties of the Haar measures and the group convolution are discussed in Appendix C.
Appendix C The $\delta$-Function of a Group
Let $G$ be a finite-dimensional Lie group with left- and right-invariant Haar measures $d_{\mathrm{L}}x\,$ and $d_{\mathrm{R}}x\,$ as defined in section IV.1.

### v2

@@ IV.3 Sequential Measurements: Kraus-Operator Distributions and the Convolution (2 of 3)
Equations 101 and 103 are deeper elaborations of equations 62 and 278.
V Kraus-Operator Densities: Ultraoperators and the IG Algebra
@@
Just like the real line or the complex plane, the IG is the domain of an entire suite of analytic tools — with invariant integrals (V.1), invariant derivatives (V.2), involutive symmetries (V.3), and intertwined differential and algebraic equations (V.4)— where functions such as the Kraus-operator density (KOD) can be calculated, propogated, and varied.
V.1 The Haar Measures: Kraus-Operator Densities and the Convolution (3 of 3)
@@
As seen in sections IV.1 and IV.2 and to be explained further in section V.4, the samples of a weak-instrument Kraus-operator distribution are the increments of a continuous invariant Markov process, increments that are locally native to another space beside the IG, or rather tangent to it. In particular, this means that it is usually easier for small times to work directly with the Kraus-operator distribution of the weak instrument and not the KOD. Therefore, it is useful to notice that the KOD of a convolution can be computed by directly applying the Kraus-operator distribution of one instrument to the KOD of the other. There are two such expressions in this regard, one for the KODs with respect to the left-invariant measure,
@@
V.2 Translation Ultraoperators and their Invariant Derivatives
@@
If a coördinate system is established, these invariant derivatives can be expanded in terms of the partial derivatives with respect to those coördinates. Such expansions can be calculated with the use of the other tool of the IMP mentioned in the introduction, the so-called MMCSD. However, this paper is intentionally focused on just one of the tools of the IMP, the KOD. For further explanation, the most explicit instance of this can be found in [2, eq. 164].
@@
V.3 The IG Algebra and Two Involutions: Gelfand-Kolmogorov and Cartan-Schmidt
@@
There are a couple of important subspaces of the IGA that come up. Although here is not the place to dig into them too much, they are mentioned because the KODs which solve the Kolmogorov equation to be introduced in section V.4 are indeed in these subspaces for finite times. The first important subspace is the closure of the space of infinitely-differentiable functions with compact support, sometimes called the Frechet space [59]. The dual space of this Frechet space is called the space of distributions. The term “distribution” here refers to “functions” like Dirac delta-functions and their derivatives which although related are not to be confused with the Kraus-operator distribution, where the KOD is actually in the Frechet space, on the other side of the function duality from the “distributions”. The second important subspace is the so-called Schwartz space for semisimple IGs, their dual space known as the tempered distributions. [60] If the KOD is in the Schwartz space, then the corresponding instrument-element representations are examples of the so-called tempered representations. [61] However, for quantum systems the instrument-element representation is generally not a unitary representation as will be explained in just a moment.
@@
which will be called the Kolmogorov (ultraoperator) adjoint because of its application in section V.4. The use of an , called the “Ankh” or “key to life”, is mostly because every obvious choice to denote an adjoint has already been taken — * for convolution, $\dagger$} for Hermitian conjugation, $\ddagger$} for the Hilbert-Schmidt adjoint (equation 268) , and $\ddagger$ for the Choi-Jamiolkowski quasi-adjoint (equation 270). As far as the author is aware, the Ankh does not have any political or historical associations that might make it a distastful symbol to use. If anything, the Ankh has a fun turn-of-the-twentieth century Western alchemical vibe.
@@
Meanwhile, the right-translation ultraoperators are not stochastic unless the IG is unimodular. We will have no further use for the right-translation ultraoperators, so this point will not be discussed further. Remember that the integration has been chosen with respect to the left-invariant Haar measure because the time of the measuring process is always considered to march from right to left, as explained at the end of section V.1.
@@
where in particular $d_{\mathrm{L}}(x^{\dagger})=d_{\mathrm{R}}x\,$ because the right-invariant measure is unique [51, p. 26]. In the same way as was explained for the , the IGA is not quite a $\text{C}^{\ddagger}$-algebra” either, but I believe it too defines another enveloping C*-algebra by the same standard construction [36, p. 47]. The enveloping $\text{C}^{\ddagger}$-algebra is clearly relevant to continuous quantum measurement. Nevertheless, it is a technical perspective that does not seem to appear yet in quantum measurement circles.
@@
V.4 The Kolmogorov Equation for the KOD of a Continuous Measurement
Let $\mathcal{I}_{dt}$ be a weak measurement such as those found in sections IV.1 and IV.2 and $d\mu_{T}(x)$ be the Kraus-operator distribution of a continuous instrument
@@ V.4 The Kolmogorov Equation for the KOD of a Continuous Measurement
Meanwhile, let the KOD — introduced in section V.1 (see equations 108 and 125) — of the instrument at time $T$ be denoted as
@@
as explained around equations 64 and 278, but so do the Kraus-operator distribution and its KOD:
@@
which can be thought of as special cases of equations 101, 103, and 114. Equations such as 191 are examples of so-called Chapman-Kolmogorov equations.[3, 64, 65, 66]
Let $\mathcal{O}_{x}$ be an instrument-element representation of the IG (as explained in section IV.3.) Observe in particular that they intertwine the left-translation ultraoperators (introduced in section V.2) with left composition by the corresponding superoperator from the instrument-element representation,
@@
it is clear that the total operation superoperator $\mathcal{Z}_{dt}$ acting on the left of the instrument-element representation $\mathcal{O}_{x}$ by composition is equivalent to the left-convolution ultraoperator (also introduced in section V.2)
@@
where is the Kolmogorov adjoint over the IG (introduced in section V.3.) Meanwhile, the adjoint of the total operation ultraoperator of the weak instrument will have an infinitesimal generator,
@@
as introduced in section IV.1 (equations 65 and 67.) Simultaneous measurements can also be considered, but for simplicity only one Lindblad operator will be considered at first. These weak instruments define IGs (also explained in section IV.1.) Let the instrument elements of these weak instruments be “minimally” separated into a Kraus-operator distribution and an instrument-element representation in the following way:
@@ VI Concluding Remarks
In addition to the Kolmogrov generator, further operators on the IGA, generally called ultraoperators, were also considered, allowing the analysis to be completely autonomous from the superoperators that originally represented the measuring process. Perhaps the most fundamental of the ultraoperators were the right-invariant derivatives and left-translations, denoted $\underleftarrow{X}$ and $\mathfrak{L}_{x}$. With these, the left-convolution ultraoperators could be constructed in direct analogy with the quantum channel superoperators,
@@
Table 4: Ultraoperator-Superoperator Intertwining Relations: The instrument-element representations of the IG, $\mathcal{O}_{x}^{(\lambda)}$ define an intertwiner between the algebra of ultraoperators (acting on the IGA) and the algebra of superoperators (acting on the dual of the von Neumann algebra). This table displays four pairs of operators, intertwined by the instrument-element representation, that are particularly relevant to continuous measurements and sequential measurements. After the IGA and ultraoperator are introduced, these intertwining relations were derived and used in section §V.4. Concepts with the ※  mark were first introduced in previous work [3].
@@
Having said all of this, it is crucial to return to the other primary tool of the IMP that was virtually abandoned at the beginning of this paper, the so-called modified Maurer-Cartan stochastic differential or MMCSD. The MMCSD is precisely what an automated measuring instrument would use to update its IG position as it continues to observe. Further, such automation is only possible if the IG is automatable — that is, itself “automatic” (section IV.1).
@@ Acknowledgements.
The author would like to thank Carl Caves, Lucien Hardy, and Rob Spekkens for their support. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Colleges and Universities.
Appendix A Quantum Measurement Theory: An Overview of Standard Knowledge
This section is a summary of all the basic tools that will be needed in this paper which should already be familiar to most quantum scientists — instruments (A.1), channels (A.2), and Lindbladians (A.3).
Section A.1 defines the generalized measurement, its Kraus operators, and the generalized Born rule. The positive-operator-valued measure (POVM) is introduced and the Luders rule is briefly mentioned. Certain operations are then considered, definining the quantum instrument and its subsequent total quantum operation. Then the all-important concept of an indirect measurement is discussed. The section finishes by addressing the position measurement.
Section A.2 then moves into the topic of decoherence channels and their usual superoperator properties, complete positivity and trace-preservation. The Hilbert-Schmidt adjoint is defined along with the Choi map and the subsequent superoperator adjoint dual to the Hilbert-Schmidt. Some care is then taken to ensure it is understood that the instrument behind the channel represents an observer and is not to be mistaken for an environment.
Section A.3 then introduces the most basic tools of continuous measurement. Lindblad operators and the Lindbladian superoperator are discussed. The one-parameter semigroup property is then explained as well as the subsequent Lindblad master equation. Although the stochastic Lindblad master equation should also be considered if we were interested in the evolution of the state during measurement, the subject is neglected because this paper is about instrument evolution.
A.1 Kraus operators, POVMs, and Instruments
$\sum_{x}\Omega_{x}^{\dagger}\Omega_{x}=1.$ (255)
$\mathcal{E}=\Big\{\Omega_{x}^{\dagger}\Omega_{x}\Big\}_{x}.$ (257)
$A\!\odot\!B^{\dagger}\,(\rho)\equiv A\rho B^{\dagger}.$ (259)
$\Omega_{x}=\sqrt{d\mu(x)}K_{x}.$ (263)
$\mathcal{E}=\{d\mu(x)K_{x}^{\dagger}K_{x}\}_{x}$ (264)
Without further context, the choice of how the Kraus-operator distribution is factored out of the instrument element is arbitrary or “ostensible” [20, 15]. In the context of continuous and sequential measurements, however, there is a distinct choice to make as will be explained in section IV.3.
$\Omega_{x}=\langle x|\mathcal{U}_{\text{int}}|0\rangle,$ (266)
A.2 Total Operations (a.k.a. Channels) and Decoherence
the Choi involution
and the Choi-Jamiolkowski quasi-adjoint
[…]
It is easy to then verify the following useful formulae
[…]
for which it is in turn easy to observe the properties
[…]
A.3 Continuous-in-Time Decoherence/Dissipation
$\mathrm{ad}[iH]\equiv-iH\odot 1+1\odot iH,$ (273)
[…]
where such an $\hat{L}$ is called a Lindblad operator [85]. The $\hat{L}^{\dagger}\hat{L}$ of the last two terms in equation 274 can be considered a non-Hermitian (see, for example [86, section 2.1.3]) perturbation to the Hamiltonian which corresponds to the shift and decay of the system’s energy levels due to interacting with an environment or meter (as described in appendix B.)
General Lindbladians are the sum of a Hamiltonian superoperator and any number of dissipators. Dissipative channels are therefore generated by infinitesimal channels of the form
$\mathcal{Z}^{(\hat{L})}_{dt}\equiv e^{\mathcal{D}[\hat{L}]\kappa dt}$ (275)
$=e^{\mathcal{D}[\hat{L}]\kappa T}$ (277)
$\mathcal{Z}_{t+\Delta t}=\mathcal{Z}_{\Delta t}\circ\mathcal{Z}_{t},$ (278)
where the semigroup can be considered to refer to the set of channels $\{\mathcal{Z}_{t}\}_{t\geq 0}$, but it often can also refer implicitly to the semigroup of composite instrument elements as will be explained later (in section IV.1, equation 68, and section IV.3, equation 103). A state that evolves according to this channel $\rho_{t}=\mathcal{Z}_{t}(\rho_{0})$ satisfies a partial differential equation
Appendix B System-Meter-Interaction Models
@@ B.2 Diffusive Processes
Appendix C Invariant One-Forms, Haar Measures, and the Modular Function
@@
Some further properties of the Haar measures and the group convolution are discussed in Appendix D.
Appendix D The $\delta$-Function of a Group
Let $G$ be a finite-dimensional Lie group with left- and right-invariant Haar measures $d_{\mathrm{L}}x\,$ and $d_{\mathrm{R}}x\,$ as defined in section V.1.